OpenStudy (anonymous):

Suppose that a polynomial function of degree 5 with rational coefficients has 2, √3, and 5+2i as zeros. Find the other zeros.

OpenStudy (anonymous):

one is 5-2i

OpenStudy (anonymous):

hint : irrational roots occur in conjugate pairs

OpenStudy (anonymous):

if \(a+\sqrt{b}\) is a zero, then \(a-\sqrt{b}\) will also be a zero

OpenStudy (anonymous):

so what does it mean @sara12345

OpenStudy (anonymous):

you're given a irrational root right ?

OpenStudy (anonymous):

since \(\sqrt{3}\) is a zero, \(-\sqrt{3}\) will also be a zero

OpenStudy (anonymous):

so how many zeroes did you get so far ?

OpenStudy (anonymous):

i really don't know...teacher just taught the lesson today...

OpenStudy (anonymous):

its okay, im asking how many zeroes *we* got so far :)

OpenStudy (anonymous):

in the question we're already given 3 zeroes : \(2, \sqrt{3} , 5+2i\) we found one more zero : \(-\sqrt{3}\)

OpenStudy (anonymous):

so total, 4 zeroes we have so far

OpenStudy (anonymous):

since its a 5 degree polynomial, there will be one more zero which we need to find okj

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

np :) you found the 5th zero also ?

OpenStudy (anonymous):

oh 5-2i

OpenStudy (anonymous):

Yes ! 5 degree polynomial will have 5 zeroes. so we're done :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np :)